5/14: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (BS, )
- 5/17~ increasing ETMY temperature.
- 5/20-24: Exchanging cryo valves.
5/2: Cooling start
- 5/29: Cooling done
- End of April~: Y arm commissioning
- ~End of May: Y arm and FPMI commissioning
Report of work
- Dew was found at ITMY, ETMY and ETMX.
- COM: Finnesse was ~28 on Thursday.
- Friday, after increasing temperature, finesse increased to ~120 with 30K.
- Sunday, temperature increased to 50K and the finesse was 140, not drastic improvement.
- PRM will turn off COACH filter on PSL for 10 min to measure acoustic back ground.
- Type B: To improve Shield for SR2 oplev
- Only ETMY temperature is increasing.
- DAQ is not stable. All connection was lost yesterday and this morning.
- IOO: It takes some time for checking MZ this morning, or afternoon if needed.
- Position of ITMX using auto leveler this afternoon or tomorrow depending work of MZ above.
- CRY: Exchanging cryo valve during this week.